Grade 12Algebra

Let s1,s2,.......and T1,T2 …......are two Arithmetic Sequences such that S1 = T1 1
T1 not 0
s2 = 2T2 and
S1 + S2 + …....S10 = T1 + T2 …...T15
Then the value of {S2 – S1} / ( T2 – T1)

Alex
7 Years agoGrade 12

1 Answer

Hanisha
7 Years ago
let first elements  and common differences in both elements are a1,a2,d1,d2
a1=a2=a due to s1= s2
a1+d1=2a2=2d2 due to s2=2t2
d1=a+2d2
4a+18d1=6a+42d2
4a+18a+36d2=6a+42d2
16a=6d2
d2=8a/3
d1=17a/3
s2-s1/t2-t1=d1/d2=17/8